surface area of 8 feet wide 8 feet high and 12 feet long?
Answer by simranb(63)

For what do you want surface area?Is it for a cuboid?
Then
S.A=2lb+2lh+2bh [l=12feet
S.A=2*12*8+2*12*8+2*8*8 b=8feet
S.A=192+192+128
S.A=512sq feet
Hope this helps!
h=8feet]